# Properties preserved by classes of Chu transforms

### v1

@@ 4. Preservation of compactness and inconsistency-flow formulas
In analogy with the preservation results of the previous section, we aim to show that Proposition 4.5 is optimal for $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formulas. First, we establish a lemma which may be of independent interest.
@@
Let $\langle X,r,A\rangle$ and $\langle Y,s,B\rangle$ be Chu spaces; given finite tuples $\bm{x}_{0}\in X$ and $\bm{y}_{0}\in Y$ of matching length, as well as finite tuples $\bm{a}_{0}\in A$ and $\bm{b}_{0}\in B$ of matching length, the following are equivalent:
@@
    for every first-order inconsistency-flow formula $\gamma(\bm{v}_{0}^{\mathsf{p}},\bm{v}_{0}^{\mathsf{s}})$, if $\langle X,r,A\rangle\models\gamma[\bm{x}_{0},\bm{a}_{0}]$ then $\langle Y,s,B\rangle\models\gamma[\bm{y}_{0},\bm{b}_{0}]$;
@@
    there exist elementary extensions $\langle X,r,A\rangle\preceq\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ and $\langle Y,s,B\rangle\preceq\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$ as well as a dense Chu transform $(f,g)$ from $\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ to $\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$ such that $f(\bm{x}_{0})=\bm{y}_{0}$ and $g(\bm{b}_{0})=\bm{a}_{0}$.
@@
  1. (i)
    $\langle X_{0},r_{0},A_{0}\rangle=\langle X,r,A\rangle$ and $\langle Y_{0},s_{0},B_{0}\rangle=\langle Y,s,B\rangle$;
  2. (ii)
    for every $n<\omega$, $\langle X_{n},r_{n},A_{n}\rangle\preceq\langle X_{n+1},r_{n+1},A_{n+1}\rangle$ and $\langle Y_{n},s_{n},B_{n}\rangle\preceq\langle Y_{n+1},s_{n+1},B_{n+1}\rangle$;
  3. (iii)
    for every $n<\omega$, $f_{n}\colon X_{n}\to Y_{n+1}$ and $g_{n}\colon B_{n+1}\to A_{n+1}$;
  4. (iv)
    $f_{0}(\bm{x}_{0})=\bm{y}_{0}$ and $g_{0}(\bm{b}_{0})=\bm{a}_{0}$;
  5. (v)
    for every $n<\omega$, $f_{n}\subseteq f_{n+1}$ and $g_{n}\subseteq g_{n+1}$;
  6. (vi)
@@
    for every $n<\omega$, every first-order inconsistency-flow formula $\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$, and parameters $\bm{x}\in X_{n}$ and $\bm{b}\in B_{n+1}$, if $\langle X_{n+1},r_{n+1},A_{n+1}\rangle\models\gamma[\bm{x},g_{n}(\bm{b})]$ then $\langle Y_{n+1},s_{n+1},B_{n+1}\rangle\models\gamma[f_{n}(\bm{x}),\bm{b}]$.
For the base step of the recursion, let $\langle X_{0},r_{0},A_{0}\rangle$ and $\langle Y_{0},s_{0},B_{0}\rangle$ be given by the first condition. We begin by constructing $\langle Y_{1},s_{1},B_{1}\rangle$ and $f_{0}\colon X_{0}\to Y_{1}$. Let $\{c_{x}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}x\in X_{0}\}$, $\{c_{y}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}y\in Y_{0}\}$, and $\{c_{b}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}b\in B_{0}\}$ be pairwise disjoint sets of new constants for elements of $X_{0}$, $Y_{0}$, and $B_{0}$ respectively. In case of a tuple $\bm{y}=\langle y_{1},\dots,y_{k}\rangle$, we write $c_{\bm{y}}=\langle c_{y_{1}},\dots,c_{y_{k}}\rangle$. Let
$\begin{split}\Delta_{0}=\bigl{\{}\varphi(c_{\bm{y}},c_{\bm{b}})\,\bigm{|}\,\mathopen{}&\varphi(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})\text{ is an }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-formula}\text{, }\bm{y}\in Y_{0},\\ &\bm{b}\in B_{0}\text{, and }\langle Y_{0},s_{0},B_{0}\rangle\models\varphi[\bm{y},\bm{b}]\bigr{\}},\end{split}$
namely the elementary diagram of $\langle Y_{0},s_{0},B_{0}\rangle$, and let
@@
$\begin{split}\Gamma_{0}=\bigl{\{}\gamma(c_{\bm{x}},c_{\bm{y}_{0}},c_{\bm{b}_{0}})\,\bigm{|}\,\mathopen{}&\gamma(\bm{v}^{\mathsf{p}},\bm{v}_{0}^{\mathsf{p}},\bm{v}_{0}^{\mathsf{s}})\text{ is first-order inconsistency-flow},\\ &\bm{x}\in X_{0}\text{, and }\langle X_{0},r_{0},A_{0}\rangle\models\gamma[\bm{x},\bm{x}_{0},\bm{a}_{0}]\bigr{\}}.\end{split}$
@@
$\exists\bm{v}_{1}^{\mathsf{p}}\dots\exists\bm{v}_{m}^{\mathsf{p}}(\gamma_{1}\land\dots\land\gamma_{m});$
then $\gamma$ is also inconsistency flow and $\langle X_{0},r_{0},A_{0}\rangle\models\gamma[\bm{x}_{0},\bm{a}_{0}]$. By (1), we have $\langle Y_{0},s_{0},B_{0}\rangle\models\gamma[\bm{y}_{0},\bm{b}_{0}]$, which means that for each $1\leq i\leq m$ there exists $\bm{y}_{i}\in Y_{0}$ such that
$\langle Y_{0},s_{0},B_{0}\rangle\models\gamma_{i}[\bm{y}_{i},\bm{y}_{0},\bm{b}_{0}].$
[…]
Let $\mathfrak{N}$ be a model of $\Delta_{0}\cup\Gamma_{0}$ and let $\langle Y_{1},s_{1},B_{1}\rangle$ be the reduct of $\mathfrak{N}$ obtained by forgetting the new constant symbols. Define $f_{0}\colon X_{0}\to Y_{1}$ as follows: for every $x\in X_{0}$, let $f_{0}(x)$ be the interpretation of the constant symbol $c_{x}$ in the structure $\mathfrak{N}$. It is clear from the construction that $\langle Y_{0},s_{0},B_{0}\rangle\preceq\langle Y_{1},s_{1},B_{1}\rangle$ and that for every first-order inconsistency-flow formula $\gamma(\bm{v}^{\mathsf{p}},\bm{v}_{0}^{\mathsf{p}},\bm{v}_{0}^{\mathsf{s}})$ and $\bm{x}\in X_{0}$,
(eq-3) $\langle X_{0},r_{0},A_{0}\rangle\models\gamma[\bm{x},\bm{x}_{0},\bm{a}_{0}]\implies\langle Y_{1},s_{1},B_{1}\rangle\models\gamma[f_{0}(\bm{x}),\bm{y}_{0},\bm{b}_{0}].$
From (eq-3) it follows already that $f_{0}(\bm{x}_{0})=\bm{y}_{0}$. Indeed, the formula $(\bm{v}^{\mathsf{p}}=\bm{v}_{0}^{\mathsf{p}})$ is inconsistency-flow and $\langle X_{0},r_{0},A_{0}\rangle\models(\bm{x}_{0}=\bm{x}_{0})$, whence $\langle Y_{1},s_{1},B_{1}\rangle\models(f_{0}(\bm{x}_{0})=\bm{y}_{0})$.
Next, we construct $\langle X_{1},r_{1},A_{1}\rangle$ and $g_{0}\colon B_{1}\to A_{1}$. Let $\{c_{x}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}x\in X_{0}\}$, $\{c_{a}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}a\in A_{0}\}$, and $\{c_{b}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}b\in B_{1}\}$ be pairwise disjoint sets of new constants for elements of $X_{0}$, $A_{0}$, and $B_{1}$ respectively. Let $\Delta_{0}^{\prime}$ be the elementary diagram of $\langle X_{0},r_{0},A_{0}\rangle$ and let
$\begin{split}\Gamma_{0}^{\prime}=\bigl{\{}\lnot\gamma(c_{\bm{x}},c_{\bm{b}},c_{\bm{a}_{0}})\,\bigm{|}\,\mathopen{}&\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}},\bm{v}_{0}^{\mathsf{s}})\text{ is first-order inconsistency-flow},\\ &\bm{x}\in X_{0}\text{, }\bm{b}\in B_{1}\text{, and }\langle Y_{1},s_{1},B_{1}\rangle\models\lnot\gamma[f_{0}(\bm{x}),\bm{b},\bm{b}_{0}]\bigr{\}}.\end{split}$
We remark that, since the class of inconsistency-flow formulas is not closed under negation, we need to consider negations of inconsistency-flow formulas in $\Gamma_{0}^{\prime}$.
We can prove that the theory $\Delta_{0}^{\prime}\cup\Gamma_{0}^{\prime}$ is consistent by an argument completely analogous to the one for $\Delta_{0}\cup\Gamma_{0}$, this time using (eq-3) and the closure of inconsistency-flow formulas under disjunctions and universal quantification over states. Let $\mathfrak{M}$ be a model of $\Delta_{0}^{\prime}\cup\Gamma_{0}^{\prime}$ and let $\langle X_{1},r_{1},A_{1}\rangle$ be the reduct of $\mathfrak{M}$. To define the function $g_{0}\colon B_{1}\to A_{1}$, for each $b\in B_{1}$ we let $g_{0}(b)$ be the interpretation of the constant symbol $c_{b}$ in the structure $\mathfrak{M}$. As before, it is clear from the construction that $\langle X_{0},r_{0},A_{0}\rangle\preceq\langle X_{1},r_{1},A_{1}\rangle$ and that for every first-order inconsistency-flow formula $\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}},\bm{v}_{0}^{\mathsf{s}})$ and parameters $\bm{x}\in X_{0}$ and $\bm{b}\in B_{1}$,
(eq-4) $\langle X_{1},r_{1},A_{1}\rangle\models\gamma[\bm{x},g_{0}(\bm{b}),\bm{a}_{0}]\implies\langle Y_{1},s_{1},B_{1}\rangle\models\gamma[f_{0}(\bm{x}),\bm{b},\bm{b_{0}}].$
Now (vi) follows from (eq-4) as a special case. Finally, using the fact that $\lnot(\bm{v}^{\mathsf{s}}=\bm{v}_{0}^{\mathsf{s}})$ is an inconsistency-flow formula, we can deduce that $g_{0}(\bm{b}_{0})=\bm{a}_{0}$, therefore point (iv) is also satisfied. This concludes the base step of the recursive construction.
Since the same method works for the general case without complications, we only sketch the details. At stage $n<\omega$, assume we have constructed Chu spaces $\langle X_{n+1},r_{n+1},A_{n+1}\rangle$ and $\langle Y_{n+1},s_{n+1},B_{n+1}\rangle$ as well as functions $f_{n}\colon X_{n}\to Y_{n+1}$ and $g_{n}\colon B_{n+1}\to A_{n+1}$ satisfying conditions (i) to (vi). To get $\langle Y_{n+2},s_{n+2},B_{n+2}\rangle$ and $f_{n+1}\colon X_{n+1}\to Y_{n+2}$, we first expand the language with new constants for elements of $X_{n+1}$, $Y_{n+1}$, and $B_{n+1}$. Then we let $\Delta_{n+1}$ be the elementary diagram of $\langle Y_{n+1},s_{n+1},B_{n+1}\rangle$ and define $\Gamma_{n+1}$ by
$\begin{split}\bigl{\{}\gamma\bigl{(}c_{\bm{x}},c_{f_{n}(\bm{x}^{\prime})},c_{\bm{b}}\bigr{)}\,\bigm{|}\,\mathopen{}&\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\prime\mathsf{p}},\bm{v}^{\mathsf{s}})\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency-flow, }\bm{x}\in X_{n+1}\text{, }\bm{x}^{\prime}\in X_{n},\\ &\bm{b}\in B_{n+1}\text{, and }\langle X_{n+1},r_{n+1},A_{n+1}\rangle\models\gamma[\bm{x},\bm{x}^{\prime},g_{n}(\bm{b})]\bigr{\}}.\end{split}$
The consistency of the theory $\Delta_{n+1}\cup\Gamma_{n+1}$ follows as before, using the inductive assumption (vi) and the closure of inconsistency-flow formulas under conjunctions and existential quantification over points. A model of $\Delta_{n+1}\cup\Gamma_{n+1}$ gives an elementary extension $\langle Y_{n+2},s_{n+2},B_{n+2}\rangle$ of $\langle Y_{n+1},s_{n+1},B_{n+1}\rangle$ and $f_{n+1}\colon X_{n+1}\to Y_{n+2}$ which satisfy the following condition: for every first-order inconsistency-flow formula $\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\prime\mathsf{p}},\bm{v}^{\mathsf{s}})$ and parameters $\bm{x}\in X_{n+1}$, $\bm{x}^{\prime}\in X_{n}$, and $\bm{b}\in B_{n+1}$,
(eq-5) $\langle X_{n+1},r_{n+1},A_{n+1}\rangle\models\gamma[\bm{x},\bm{x}^{\prime},g_{n}(\bm{b})]\implies\langle Y_{n+2},s_{n+2},B_{n+2}\rangle\models\gamma[f_{n+1}(\bm{x}),f_{n}(\bm{x}^{\prime}),\bm{b}].$
Since the formula $(v^{\mathsf{p}}=v^{\prime\mathsf{p}})$ is inconsistency-flow, (eq-5) implies in particular that $f_{n+1}(x^{\prime})=f_{n}(x^{\prime})$ for each $x^{\prime}\in X_{n}$, whence $f_{n}\subseteq f_{n+1}$.
Assuming (eq-5), we can build $\langle X_{n+2},r_{n+2},A_{n+2}\rangle$ and $g_{n+1}\colon B_{n+2}\to A_{n+2}$ using the same method, thus completing the recursive construction.
Now define
$\langle X^{\prime},r^{\prime},A^{\prime}\rangle=\biggl{\langle}\bigcup_{n<\omega}X_{n},\bigcup_{n<\omega}r_{n},\bigcup_{n<\omega}A_{n}\biggr{\rangle}$
and
$\langle Y^{\prime},s^{\prime},B^{\prime}\rangle=\biggl{\langle}\bigcup_{n<\omega}Y_{n},\bigcup_{n<\omega}s_{n},\bigcup_{n<\omega}B_{n}\biggr{\rangle}.$
By points (i) and (ii) and the elementary chain theorem of Alfred Tarski and Robert Vaught [30, Theorem 1.9], we have $\langle X,r,A\rangle\preceq\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ and $\langle Y,s,B\rangle\preceq\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$.
Let $f=\bigcup_{n<\omega}f_{n}$ and $g=\bigcup_{n<\omega}g_{n}$. By points (iv) and (v), this gives indeed two functions $f\colon X^{\prime}\to Y^{\prime}$ and $g\colon B^{\prime}\to A^{\prime}$ such that $f(\bm{x}_{0})=\bm{y}_{0}$ and $g(\bm{b}_{0})=\bm{a}_{0}$.
Finally, we use (vi) to verify that $(f,g)$ is a dense Chu transform from $\langle X^{\prime},r^{\prime},A^{\prime}\rangle$ to $\langle Y^{\prime},s^{\prime},B^{\prime}\rangle$. First, to check the adjointness condition of Definition 2.6, take $x\in X^{\prime}$ and $b\in B^{\prime}$ and choose $n<\omega$ sufficiently large that $x\in X_{n}$ and $b\in B_{n+1}$. Since both $R(v^{\mathsf{p}},v^{\mathsf{s}})$ and $\lnot R(v^{\mathsf{p}},v^{\mathsf{s}})$ are inconsistency-flow formulas, point (vi) implies that
$\langle x,g(b)\rangle\in r^{\prime}\iff\langle x,g_{n}(b)\rangle\in r_{n+1}\iff\langle f_{n}(x),b\rangle\in s_{n+1}\iff\langle f(x),b\rangle\in s^{\prime}.$
For the density condition, let $b\in B^{\prime}$ and suppose there exists $y\in Y^{\prime}$ such that $\langle y,b\rangle\in s^{\prime}$. Let $n<\omega$ be sufficiently large that $\langle y,b\rangle\in s_{n+1}$, hence in particular $\langle Y_{n+1},s_{n+1},B_{n+1}\rangle\models\exists v^{\mathsf{p}}R(v^{\mathsf{p}},b)$. Since $\lnot\exists v^{\mathsf{p}}R(v^{\mathsf{p}},v^{\mathsf{s}})$ is an inconsistency-flow formula, by (vi) it follows that $\langle X_{n+1},r_{n+1},A_{n+1}\rangle\models\exists v^{\mathsf{p}}R(v^{\mathsf{p}},g_{n}(b))$. Choose $x\in X_{n+1}$ such that $\langle x,g_{n}(b)\rangle\in r_{n+1}$; then $\langle x,g(b)\rangle\in r^{\prime}$ and finally $\langle f(x),b\rangle\in s^{\prime}$, as we had to show.
[…]
@@
Chen Chung Chang and H. Jerome Keisler [8, Section 3.2] applied the method of elementary chains to obtain several classic preservation results. In a similar fashion, we can apply Lemma 4.11 to obtain our main theorem.

### v3

@@ 4. Preservation of compactness and inconsistency-flow formulas
In analogy with the preservation results of the previous section, we aim to show that Proposition 4.5 is optimal for $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formulas. To do so, we shall use recursively saturated model pairs as in van Benthem’s proof of Theorem 3.6. Recursively saturated structures were first studied in John Schlipf’s PhD thesis [33] and later appeared in a paper by Barwise and Schlipf [7]. Let us first recall their definition.
Definition 4.11 (Barwise and Schlipf [7, Definition 1.1]).
Let $\mathcal{L}$ be a finite language. An $\mathcal{L}$-structure $\mathfrak{M}$ is recursively saturated if for every finite $C\subseteq M$ and every recursive set $\Gamma(x)$ of $\mathcal{L}_{\omega,\omega}$-formulas with parameters from $C$, if $\Gamma(x)$ is finitely satisfiable in $\mathfrak{M}$, then $\Gamma(x)$ is satisfiable in $\mathfrak{M}$.
The next theorem is an existence result for recursively saturated structures.
Theorem 4.12 (Barwise and Schlipf [7, 1.4(i)]).
Each structure in a finite language has a recursively saturated elementary extension of the same cardinality.
Let $\mathcal{L}=\{R_{1},\dots,R_{n}\}$ be a finite relational language. Given two $\mathcal{L}$-structures $\mathfrak{M}$ and $\mathfrak{N}$, the model pair $(\mathfrak{M},\mathfrak{N})$ is defined as the structure
$(\mathfrak{M},\mathfrak{N})=\bigl\langle M\cup N,M,N,R_{1}^{\mathfrak{M}},\dots,R_{n}^{\mathfrak{M}},R_{1}^{\mathfrak{N}},\dots,R_{n}^{\mathfrak{N}}\bigr\rangle.$
Each $\mathcal{L}_{\omega,\omega}$-formula $\varphi(\bm{x})$ has relativizations $\varphi^{\prime}(\bm{x})$ and $\varphi^{\prime\prime}(\bm{x})$, obtained from $\varphi$ in a recursive way, such that for all $\bm{m}\in M$
$\mathfrak{M}\models\varphi[\bm{m}]\iff(\mathfrak{M},\mathfrak{N})\models\varphi^{\prime}[\bm{m}]$
and for all $\bm{n}\in N$
$\mathfrak{N}\models\varphi[\bm{n}]\iff(\mathfrak{M},\mathfrak{N})\models\varphi^{\prime\prime}[\bm{n}].$
This technique of gluing structures together dates back to Abraham Robinson’s proof of his consistency theorem [29, Theorem 2.6]. In the next lemma, we apply this to Chu spaces coded as two-sorted relational structures.
@@
Let $\langle X,r,A\rangle$ and $\langle Y,s,B\rangle$ be countable Chu spaces such that the model pair $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$ is recursively saturated. Given finite tuples $\bm{x}\in X$ and $\bm{y}\in Y$ of matching length, as well as finite tuples $\bm{a}\in A$ and $\bm{b}\in B$ of matching length, the following are equivalent:
@@
    for every first-order inconsistency-flow formula $\gamma(\bm{v}^{\mathsf{p}},\bm{v}^{\mathsf{s}})$, if $\langle X,r,A\rangle\models\gamma[\bm{x},\bm{a}]$ then $\langle Y,s,B\rangle\models\gamma[\bm{y},\bm{b}]$;
@@
    there exists a dense Chu transform $(f,g)$ from $\langle X,r,A\rangle$ to $\langle Y,s,B\rangle$ such that $f(\bm{x})=\bm{y}$ and $g(\bm{b})=\bm{a}$.
@@
$X=\{x_{n}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}n<\omega\}\quad\text{and}\quad B=\{b_{n}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}n<\omega\}$
in such a way that $x_{0}=x$ and $b_{0}=b$.
We shall recursively construct $\langle y_{n}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}n<\omega\rangle$ in $Y$ and $\langle a_{n}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}n<\omega\rangle$ in $A$ such that for each $n<\omega$ the following condition is satisfied:
@@
(3) $\text{for every first-order inconsistency-flow formula }\gamma(v_{0}^{\mathsf{p}},\dots,v_{n}^{\mathsf{p}},v_{0}^{\mathsf{s}},\dots,v_{n}^{\mathsf{s}}),\\ \langle X,r,A\rangle\models\gamma[x_{0},\dots,x_{n},a_{0},\dots,a_{n}]\implies\langle Y,s,B\rangle\models\gamma[y_{0},\dots,y_{n},b_{0},\dots,b_{n}].$
For the base case, let $y_{0}=y$ and $a_{0}=a$. Condition (3) is satisfied because of our assumption (1).
At stage $n<\omega$, assume we have constructed $\langle y_{0},\dots,y_{n}\rangle$ and $\langle a_{0},\dots,a_{n}\rangle$ satisfying (3). To get $y_{n+1}\in Y$, consider the following recursive [^5] The set of all first-order inconsistency-flow formulas is recursive. Notice that $\Gamma(v^{\mathsf{p}})$ contains no mention of the relation of satisfiability in $\mathfrak{M}$ or $\mathfrak{N}$. set of formulas:
$\Gamma(v^{\mathsf{p}})=\bigl\{\gamma^{\prime}(x_{0},\dots,x_{n+1},a_{0},\dots,a_{n})\rightarrow\gamma^{\prime\prime}(y_{0},\dots,y_{n},v^{\mathsf{p}},b_{0},\dots,b_{n})\bigm|\\ \gamma(v_{0}^{\mathsf{p}},\dots,v_{n}^{\mathsf{p}},v^{\mathsf{p}},v_{0}^{\mathsf{s}},\dots,v_{n}^{\mathsf{s}})\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency flow}\bigr\}.$
We claim that $\Gamma(v^{\mathsf{p}})$ is finitely satisfiable in $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$. Indeed, let $\gamma_{1},\dots,\gamma_{m}$ for some $m<\omega$ be inconsistency-flow formulas, as in the definition of $\Gamma(v^{\mathsf{p}})$, and suppose that for each $1\leq i\leq m$
$\langle X,r,A\rangle\models\gamma_{i}[x_{0},\dots,x_{n+1},a_{0},\dots,a_{n}].$
We let $\gamma(v_{0}^{\mathsf{p}},\dots,v_{n}^{\mathsf{p}},v_{0}^{\mathsf{s}},\dots,v_{n}^{\mathsf{s}})$ be the formula $\exists v^{\mathsf{p}}(\gamma_{1}\land\dots\land\gamma_{m})$. Since $\gamma$ is also inconsistency flow and $\langle X,r,A\rangle\models\gamma[x_{0},\dots,x_{n},a_{0},\dots,a_{n}]$, the inductive hypothesis (3) guarantees that $\langle Y,s,B\rangle\models\gamma[y_{0},\dots,y_{n},b_{0},\dots,b_{n}]$. This means that there exists $y\in Y$ such that for each $1\leq i\leq m$
$\langle Y,s,B\rangle\models\gamma_{i}[y_{0},\dots,y_{n},y,b_{0},\dots,b_{n}],$
thus showing that $\Gamma(v^{\mathsf{p}})$ is finitely satisfiable in $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$. By recursive saturation, $\Gamma(v^{\mathsf{p}})$ is satisfiable in $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$ by some element $y_{n+1}\in Y$. Then it is clear by construction that
@@
(4) $\text{for every first-order inconsistency-flow formula }\gamma(v_{0}^{\mathsf{p}},\dots,v_{n+1}^{\mathsf{p}},v_{0}^{\mathsf{s}},\dots,v_{n}^{\mathsf{s}}),\\ \langle X,r,A\rangle\models\gamma[x_{0},\dots,x_{n+1},a_{0},\dots,a_{n}]\implies\langle Y,s,B\rangle\models\gamma[y_{0},\dots,y_{n+1},b_{0},\dots,b_{n}].$
To get $a_{n+1}\in A$, we proceed analogously and consider the recursive set of formulas
$\Delta(v^{\mathsf{s}})=\bigl\{\gamma^{\prime}(x_{0},\dots,x_{n+1},a_{0},\dots,a_{n},v^{\mathsf{s}})\rightarrow\gamma^{\prime\prime}(y_{0},\dots,y_{n+1},b_{0},\dots,b_{n+1})\bigm|\\ \gamma(v_{0}^{\mathsf{p}},\dots,v_{n+1}^{\mathsf{p}},v_{0}^{\mathsf{s}},\dots,v_{n}^{\mathsf{s}},v^{\mathsf{s}})\text{ is }\mathcal{L}^{\mathrm{II}}_{\omega,\omega}\text{-inconsistency flow}\bigr\}.$
To see that $\Delta(v^{\mathsf{s}})$ is finitely satisfiable in $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$, let $\gamma_{1},\dots,\gamma_{m}$ for some $m<\omega$ be inconsistency-flow formulas as above and suppose that for each $1\leq i\leq m$
$\langle Y,s,B\rangle\models\lnot\gamma_{i}[y_{0},\dots,y_{n+1},b_{0},\dots,b_{n+1}].$
This time, we let $\gamma(v_{0}^{\mathsf{p}},\dots,v_{n+1}^{\mathsf{p}},v_{0}^{\mathsf{s}},\dots,v_{n}^{\mathsf{s}})$ be the inconsistency-flow formula $\forall v^{\mathsf{s}}(\gamma_{1}\lor\dots\lor\gamma_{m})$. Since $\langle Y,s,B\rangle\models\lnot\gamma[y_{0},\dots,y_{n+1},b_{0},\dots,b_{n}]$, property (4) implies that $\langle X,r,A\rangle\models\lnot\gamma[x_{0},\dots,x_{n+1},a_{0},\dots,a_{n}]$. This means that there exists $a\in A$ such that for each $1\leq i\leq m$
$\langle X,r,A\rangle\models\lnot\gamma_{i}[x_{0},\dots,x_{n+1},a_{0},\dots,a_{n},a],$
thus showing that $\Delta(v^{\mathsf{s}})$ is finitely satisfiable in $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$. By recursive saturation, we can pick $a_{n+1}\in A$ which satisfies $\Delta(v^{\mathsf{s}})$ in $(\langle X,r,A\rangle,\langle Y,s,B\rangle)$. This completes the recursive construction of $\langle y_{n}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}n<\omega\rangle$ and $\langle a_{n}\mathchoice{\>}{\>}{\,}{\,}|\allowbreak\mathchoice{\>}{\>}{\,}{\,}\mathopen{}n<\omega\rangle$.
Now define $f\colon X\to Y$ and $g\colon B\to A$ by letting $f(x_{n})=y_{n}$ and $g(b_{n})=a_{n}$ for each $n<\omega$. Note that, if $x_{m}=x_{n}$ for some $m,n<\omega$, then $y_{m}=y_{n}$ by (3) and the fact that $(v_{0}^{\mathsf{p}}=v_{1}^{\mathsf{p}})$ is an inconsistency-flow formula. Similarly, if $b_{m}=b_{n}$ for some $m,n<\omega$, then $a_{m}=a_{n}$ by (3) again and the fact that $\lnot(v^{\mathsf{s}}_{0}=v^{\mathsf{s}}_{1})$ is an inconsistency-flow formula. Hence, the functions $f$ and $g$ are well defined and, by the base step of the recursive construction, they satisfy $f(x)=y$ and $g(b)=a$.
Finally, we use (3) to verify that $(f,g)$ is a dense Chu transform from $\langle X,r,A\rangle$ to $\langle Y,s,B\rangle$. First, to check the adjointness condition of Definition 2.6, take $x\in X$ and $b\in B$ and choose $m,n<\omega$ such that $x_{m}=x$ and $b_{n}=b$. Since both $R(v^{\mathsf{p}},v^{\mathsf{s}})$ and $\lnot R(v^{\mathsf{p}},v^{\mathsf{s}})$ are inconsistency-flow formulas, condition (3) implies that
$\begin{split}\langle x,g(b)\rangle\in r&\iff\langle X,r,A\rangle\models R(x_{m},a_{n})\\ &\iff\langle Y,s,B\rangle\models R(y_{m},b_{n})\\ &\iff\langle f(x),b\rangle\in s.\end{split}$
For the density condition, let $b\in B$ and suppose there exists $y\in Y$ such that $\langle y,b\rangle\in s$. Choose $n<\omega$ such that $b_{n}=b$, hence in particular $\langle Y,s,B\rangle\models\exists v^{\mathsf{p}}R(v^{\mathsf{p}},b_{n})$. Since $\lnot\exists v^{\mathsf{p}}R(v^{\mathsf{p}},v^{\mathsf{s}})$ is an inconsistency-flow formula, from (3) it follows that $\langle X,r,A\rangle\models\exists v^{\mathsf{p}}R(v^{\mathsf{p}},a_{n})$. Choose $x\in X$ such that $\langle x,a_{n}\rangle\in r$; then $\langle f(x),b\rangle\in s$, as we had to show.
@@
H. Jerome Keisler [19, §5] applied the method of recursively saturated model pairs to obtain classic preservation results. In a similar fashion, we can apply Lemma 4.13 to obtain our main theorem.